OpenStudy (anonymous):

Find the variance and standard deviation. {3, 3, 4, 5, 5}

OpenStudy (anonymous):

I know the mean is 4 if that helps any in the problem

OpenStudy (kirbykirby):

So the standard deviation is found using the formula: \[ \large s=\sqrt{\frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2}\]

OpenStudy (kirbykirby):

the variance is simply the square of that, so \[ \large s^2=\frac{1}{n-1}\sum_{i=1}^n(x_i-\bar{x})^2\]

OpenStudy (anonymous):

Okay thanks, so what would i plug in in those variables?

OpenStudy (kirbykirby):

\(x_i\) are the individual data points, and \(\bar{x}\) is the mean. \(n\) is the number of data points in your set (there are 5 in your set ). \[\large s^2=\frac{1}{5-1}\left( (3-4)^2+(3-4)^4+(4-4)^2+(5-4)^2+(5-4)^2 \right) \]

OpenStudy (kirbykirby):

you're subtracting each data point with the mean, 4, then squaring that result. Then you add up of those of results together (I just realized I made a typo on the second sum above... it should be \(\large (3-4)^2\), not \(\large (3-4)^4\)).